# Neural Networks for 3D Characterization of AGATA Crystals

### v1

@@ Abstract
Precise localization of gamma-ray interactions is crucial for the performance of AGATA. The Pulse Shape Analysis (PSA) method used for the position estimation relies on simulated signals database. The Pulse Shape Comparison Scanning (PSCS) method was used to scan AGATA crystals in order to produce an experimental database. This paper presents a novel approach using LSTM neural networks to determine the 3D interaction position of gamma rays within AGATA crystals, trained on data from the Strasbourg scanning tables, allowing for constructing experimental database. A custom masked loss function is introduced to enable training with incomplete position information. The database generated by this new method outperforms the simulated database used so far, and the experimental database obtained from the conventional PSCS algorithm. Crucially, the neural network approach effectively generalizes even when trained with only one known coordinate per interaction, paving the way for simplified 1D scanning procedures.
@@
AGATA, HPGe Detectors, Pulse Shape Analysis, Gamma-Ray Tracking, Machine Learning, LSTM, Position Reconstruction, Detector Characterization
@@ 1 Introduction
Gamma-ray spectroscopy is a fundamental technique in nuclear physics, providing insights into nuclear structure and reactions. The Advanced GAmma Tracking Array (AGATA) [1, 2] represents a significant advancement in this field, particularly for in-beam experiments, utilizing highly segmented high-purity germanium (HPGe) crystals to achieve state-of-the-art energy and position resolution. The precise determination of gamma-ray interaction positions within the detector volume is essential for effective gamma-ray tracking and Doppler correction, ultimately influencing the overall performance of the array.
@@
Current methods for interaction localization rely on Pulse Shape Analysis (PSA) [3, 4, 5]. PSA algorithms compare measured signals (traces) from the detector with a pre-calculated basis of signals corresponding to known positions. These bases can be generated through simulations, such as those provided by the AGATA Detector Library (ADL) [6], or more recently using the AGATAGeFEM package [7], which utilizes a finite-element method.
@@
Several techniques have been developed within the collaboration to study the detector response and create experimental bases, called scanning tables. These tables can be divided into three techniques: coincidence scanning [8, 9], Pulse Shape Comparison Scanning (PSCS) with collimated $\gamma$ source [10], and PSCS with electronic collimation [11, 12].
@@
The PSCS with collimated source was first implemented at the Institut pluridisciplinaire Hubert Curien (IPHC) [13] (Strasbourg scanning table). In this technique, two scans are performed: one with the detector positioned horizontally relative to a collimated source and one positioned vertically. The position of the collimator constrains the position of the $\gamma$-ray interaction in the detector. Horizontal and vertical scans are therefore used to determine the ($X,Z$) and ($X,Y$) positions of the interaction, respectively. The signals from both scans are then compared to build a 3D basis of signals using the PSCS algorithm [13, 14, 15, 16] which is computationally intensive and time-consuming.
@@
Machine learning (ML) techniques, particularly neural networks (NN), have emerged as powerful tools for data analysis and pattern recognition in various fields, including nuclear physics [17]. Recent applications of ML in nuclear instrumentation include particle tracking, event classification, and signal processing [17]. The ability of NNs to learn complex, non-linear relationships and patterns directly from the data makes them ideally suited to address the challenges of gamma-ray interaction localization using detector signals.
@@
Figure 1 presents an example of super-traces corresponding to simulated gamma-ray interactions using the ADL simulation code. The figure depicts five interactions that occur at different radial positions (R) within a single segment, designated as the fired segment. The AGATA crystal nomenclature, also shown, identifies each of the 36 segments by a sector (A-F) and a layer (1-6). The hexagonal diagram (top right) illustrates the six angular sectors, while the 3D crystal model (bottom right) shows the six layers along the z-axis. In this example, A3 is the fired segment, referring to the segment in Sector A and Layer 3. The plots display the primary, net-charge signals from the central Core electrode and the fired segment (A3), alongside the transient (or induced) signals from neighboring segments (B3, F3, A2, A4). The shapes, amplitudes, and even polarities of these transient signals are highly dependent on the interaction’s radial position. For instance, the transient signal in the outer neighbouring segment (A4) changes from negative to positive as the interaction radius decreases. It is this position-dependent response across the entire super-trace that forms the basis for both the PSA and the pattern recognition performed by neural networks.
@@
Figure 1: Example of ADL simulated AGATA detector signals (super-traces) for interactions at different radial positions (R) within segment A3. The AGATA nomenclature (top right) and crystal structure (bottom right) are shown for reference. The varying shapes and polarities of the transient signals in neighbouring segments (B3, F3, A2, A4) highlight the position sensitivity used for PSA and NN-based positioning.
@@ 2 Methodology
Two AGATA crystals were scanned, S001 [14] and A005 using the Strasbourg scanning table. The S001 is a non-standard, prototype, symmetric AGATA crystal, while the A005 is a standard AGATA crystal (type A) which had been used in the AGATA campaign at GANIL [18]. More details on the AGATA crystals geometries can be found in references [1, 19].
Two 2D scans were performed for each crystal using a 137Cs source with the 1 mm diameter pinhole collimator. One with the crystal placed horizontally and one vertically. For the S001 scan, using a $2\times 2$ mm2 pitch, the vertical scan yielded 1351 points and the horizontal scan yielded 2195 points. For the A005 scan, the same pitch resulted in 1305 points for the vertical scan and 1858 points for the horizontal scan.
@@ 2.1 PSCS
The PSCS algorithm is a traditional technique for generating experimental bases from scanning table data. The PSCS algorithm relies on comparing signals from two intersecting scans (horizontal and vertical). The underlying principle is that signals originating from the same physical interaction point should have similar pulse shapes, regardless of the scan direction. A $\chi^{2}$ minimization is employed to quantify the similarity between signals from the two scans using the following formula:
@@
where $V_{ch}(i)$ and $H_{ch}(i)$ represent the signal amplitudes of channel $ch$ at time sample $i$ (where each sample is 10 ns) for the vertical and horizontal scans, respectively; $\sigma$ is the noise amplitude; and $N$ is the total number of samples compared (37 channels × 120 samples). Signal pairs with a $\chi^{2}$ value below a predefined threshold are considered to originate from the intersection point of the two scans. This process is iterative, refining the selection of signals to create a basis representative of the interaction position. However, the optimal $\chi^{2}$ threshold varies depending on the location within the crystal, requiring careful adjustment and complicating the automation of the process. This, combined with the inherently large number of comparisons required for each position, leads to significant computational cost and complexity in the implementation of the PSCS algorithm.
PSCS-derived bases for the S001 crystal [14] were provided by collaborators at IPHC Strasbourg for comparison with the neural network-generated bases.
@@ 2.2 Data Preprocessing
The raw data acquired from the TNT2 [20] card during the scans underwent several pre-processing steps:
@@
    Conversion and Compression: Raw data was converted to the AGATA Data Format (ADF) [21] using the Scanning Table Data Processing (STDPro) software [22]. This step also compressed the data significantly ($\sim$ factor of 5).
@@
    Coincidence Filtering: To filter out random coincidences, events occurring in segments with low statistics were removed. A baseline was established by calculating the average event count from all segments not illuminated by the source. Any event whose fired segment had a total count less than three times this baseline was subsequently filtered from the dataset.
@@
    Timing Alignment: Core and segment traces were time-aligned using standard AGATA workflow procedures.
@@
    Trace Sampling: Super-traces were constructed. From the original 120 samples per trace, the 60 samples corresponding to the rising edge and plateau were retained, resulting in a super-trace dimension of $(36\text{ segments}+1\text{ core})\times 60\text{ samples}=37\times 60=2220$ features per event.
@@
    Format Conversion: Data was converted from ROOT format to NumPy arrays using the ‘uproot‘ library [23] for compatibility with machine learning frameworks.
@@
    Dataset Preparation: Data from both vertical and horizontal scans for a given crystal were combined and shuffled. The combined dataset was saved in HDF5 format which allows for loading specific events from the files without the need to load the entire file into memory, a feature that becomes extremely important when building the experimental basis.
@@
    Train/Validation Split: The full dataset for each crystal ($10^{8}$ events for S001, and $0.7\times 10^{8}$ events for A005) was split into a training set (90%) and a validation set (10%) to monitor model performance and prevent overfitting.
The training dataset for the NN model consisted of events within the energy ranges of 300-480 keV (upper Compton continuum) and 650-670 keV (photopeak region). This selection was based on preliminary analysis and the findings illustrated in Fig. 7, which indicate that these energy windows correspond to the lowest prediction errors achievable by the model. By concentrating the training on these high-fidelity regions, the aim was to optimize the model’s ability to learn the position-dependent features from the most informative signals, while reducing the impact of potentially less reliable data from lower energies or complex scattering regions known to exhibit higher errors and potentially more background contamination.
@@ 2.3.2 Network Architecture
The NN architecture employed is based on Long Short-Term Memory (LSTM) layers [24]. As the detector signals (super-traces) represent time-sequential data where the temporal evolution encodes position information, LSTMs are particularly suitable. They are a type of Recurrent Neural Network (RNN) specifically designed to effectively learn and remember long-term dependencies within sequences by overcoming issues like the vanishing gradient problem that can hinder simpler RNNs when processing long sequences. This capability is crucial for capturing the subtle features across the entire pulse shape relevant for accurate position determination. Moreover, LSTMs has been shown to be robust to time misalignment [25], which is crucial to process AGATA signals. These LSTM layers are followed by dense layers for the final position regression. The input to the model is the flattened super-trace vector of 2220 nodes. The detailed architecture, illustrated schematically in Fig. 2, consists of the following layers:
@@
Figure 2: Schematic illustrating the NN architecture.
An ablation study indicated that increasing the number or size of dense layers did not improve accuracy, while the number of LSTM units affected convergence speed but not final accuracy if training iterations were adjusted proportionally.
@@ 2.3.3 Training Paradigm
The models were trained using the Adam optimizer [26] with a learning rate of $10^{-4}$. A batch size of 500 super-traces per iteration was used, with a total of $3\times 10^{4}$ iterations per model. The super-traces were normalized by scaling each one so that its peak amplitude equaled 1000 arbitrary units. This normalization procedure relies on the standard assumption that the fundamental shape of a signal is independent of the total energy deposited. This approach is consistent with the established AGATA PSA methodology. The masked Euclidean distance (Eq. 2) served as the primary loss function. Throughout this work, the ”error” is defined as the loss value, unless otherwise specified.
@@
  • Training on a uniformly sampled subset (equal number of super-traces per scanned point): This drastically reduced the dataset size (to  2% of original) and resulted in poorer overall performance [27].
@@
As shown in Fig. 3, training per segment yielded the lowest average error across segments and mitigated the distribution bias observed with a single global model. All subsequent results are based on this per-segment training strategy.
Figure 3: Error ($m$) on the validation dataset at photopeak energy versus segment number for different training methods: using a uniformly sampled subset (blue), using the full dataset with a single model trained with balanced batches per segment (orange), and training one model per segment (green). Error is calculated on the two known axes. The segments are labeled with a numerical index from 0 to 35. This index correlates directly with the standard AGATA nomenclature, where segments 0-5 correspond to Layer 1 (Sectors A-F), segments 6-11 correspond to Layer 2 (Sectors A-F), and so on, up to segments 30-35 for Layer 6.
@@ 2.3.4 Overfitting Mitigation
Despite the masked loss function, signs of overfitting were observed: the training loss continued to decrease while the validation loss reached a minimum and started to increase (as seen in Fig. 4, blue and orange curves, respectively).
To combat this overfitting and encourage generalization, a Gaussian noise layer was added after the input layer in the network architecture. This layer adds random noise, drawn from a Gaussian distribution, to each sample of the input super-trace during training. Various noise levels (standard deviations) were experimentally tested during model development. While higher noise levels could further reduce the discrepancy between training and validation loss, they also tended to degrade the overall prediction accuracy (increase the minimum validation loss). Conversely, lower noise levels were less effective at controlling the overfitting. A standard deviation of 5 keV was ultimately selected as it provided the best observed balance between these competing factors. As shown in Fig. 4, the addition of this 5 keV noise significantly reduced the gap between training and validation loss (green and red curves, respectively) compared to training without noise, and more importantly, led to a lower converged validation loss, indicating improved model generalization.

### v2

# Neural Networks for 3D Characterisation of AGATA Crystals

@@ Abstract
Precise localisation of gamma-ray interactions is crucial for the performance of the Advanced GAmma Tracking Array (AGATA). The Pulse Shape Analysis (PSA) method used for the position estimation of gamma-ray interactions relies on a simulated signal database. The Pulse Shape Comparison Scanning (PSCS) method was used to scan AGATA crystals in order to produce an experimental database of signals. This paper presents a novel approach using Long Short-Term Memory (LSTM) neural networks to determine the 3D interaction position of gamma rays within AGATA crystals, trained on data from IPHC Strasbourg, allowing for the construction of an experimental database. A custom masked loss function is introduced to enable training with incomplete position information. The database generated by this new method outperforms the existing simulated database, and the experimental database obtained from the conventional PSCS algorithm.
@@
AGATA, HPGe Detectors, Pulse Shape Analysis, Gamma-Ray Tracking, Machine Learning, LSTM, Position Reconstruction, Detector Characterisation
@@ 1 Introduction
Gamma-ray spectroscopy serves as a powerful technique in experimental nuclear physics, providing insights into nuclear structure and reactions. The Advanced GAmma Tracking Array (AGATA) [Akkoyun2012, AGATA2] significantly enhances in-beam gamma-ray spectroscopy capabilities. Through the use of highly segmented high-purity germanium (HPGe) detectors, it delivers state-of-the-art energy and position resolution. The precise determination of gamma-ray interaction positions within the detector volume is essential for effective gamma-ray tracking and Doppler correction, ultimately influencing the overall performance of the array.
@@
Current methods for interaction localisation rely on Pulse Shape Analysis (PSA) [PSA2, PSA3, PSA]. PSA algorithms compare measured signals (traces) from the detector with a pre-calculated basis of signals corresponding to known positions. These bases can be generated through simulations, such as those provided by the AGATA Detector Library (ADL) [ADL], or more recently using the AGATAGeFEM package [AGATAGeFEM].
@@
Several techniques have been developed within the collaboration to study the detector response and create experimental bases, called scanning tables. These tables can be divided into three techniques: coincidence scanning [nelson2007characterisation, ha2013new], Pulse Shape Comparison Scanning (PSCS) with a collimated $\gamma$ source [PSCS0], and PSCS with electronic collimation [goel2011spatial, hernandez2013characterization].
@@
The PSCS with a collimated source was first implemented at the Institut pluridisciplinaire Hubert Curien (IPHC) [PSCS] (Strasbourg scanning table). In this technique, two scans are performed: one with the detector positioned horizontally relative to a collimated source and one positioned vertically. The position of the collimator constrains the position of the $\gamma$-ray interaction in the detector. Horizontal and vertical scans are therefore used to determine the ($X,Z$) and ($X,Y$) positions of the interaction, respectively. The signals from both scans are then compared to build a 3D basis of signals using the PSCS algorithm [PSCS, decan2020, PSCS2, PSCS3]. A limitation of this approach is that it is computationally intensive and time-consuming, taking approximately $4.7$ days (on a $2.60$ GHz processor machine) to process data from one scanned crystal [PSCS2].
@@
Machine learning (ML) techniques, particularly neural networks (NN), have emerged as powerful tools for data analysis and pattern recognition in various fields, including nuclear physics [ML_NP]. Recent applications of ML in nuclear instrumentation include particle tracking, event classification, and signal processing [ML_NP]. The ability of NNs to learn complex, non-linear relationships and patterns directly from the data makes them ideally suited to address the challenges of gamma-ray interaction localisation using detector signals.
The AGATA spectrometer is designed to comprise tapered hexagonal coaxial n-type HPGe crystals, arranged to cover a solid angle of $4\pi$. To achieve this spherical tessellation, the array utilises three distinct asymmetric crystal shapes, designated as types A, B, and C. Despite the slight variations in shape, all crystals share a common length of 90 mm and a rear diameter of 80 mm.
To enable the precise tracking required for such a setup, each crystal is electrically segmented in 6 angular sectors and 6 depth layers, partitioning the volume into 36 segments. This outer segmentation is complemented by a central inner electrode that serves as the common anode, referred to as the core. The nomenclature used to identify these segments is illustrated in Fig. 1. The hexagonal diagram (top right) shows the six angular sectors (A-F), while the 3D crystal model (bottom right) depicts the six depth layers (1-6) along the z-axis.
@@
Figure 1 illustrates the detector response to simulated gamma-ray interactions using the ADL code. In this context, the signal from a single electrode once treated and filtered is referred to as a trace, while the complete set of signals from all 36 segments and the central contact (core trace) is collectively termed a super-trace. The figure depicts specific traces for five interactions that occur at different radial positions (R) within a single segment, designated as the fired segment. In this example, A3 is the fired segment, referring to the segment in Sector A and Layer 3. The plots display the primary net-charge traces of the core and the fired segment (A3), alongside the transient (or induced) traces from the neighbouring segments (B3, F3, A2, A4). The shapes, amplitudes, and polarities of these transient traces are highly dependent on the interaction’s radial and azimuthal position. For example, the transient trace in the outer neighbouring segment (A4) changes from negative to positive as the interaction radius decreases. It is this position-dependent response across the super-trace that forms the basis for position determination.
@@
Figure 1: An example of ADL simulated AGATA detector signals (traces) for interactions at different radial positions (R) within segment A3. The AGATA nomenclature (top right) and crystal structure (bottom right) are shown for reference. The varying shapes and polarities of the transient signals in neighbouring segments (B3, F3, A2, A4) highlight the position sensitivity used for PSA and NN-based positioning.
@@ 2 Methodology
Two AGATA crystals were characterised, S001 [decan2020] and A005 using the Strasbourg scanning table. The S001 is a prototype, symmetric AGATA crystal, while the A005 is an asymmetric AGATA crystal (type A) which had been used in the AGATA campaign at GANIL [GANIL_AGATA]. More details on the AGATA crystals geometries can be found in references [Akkoyun2012, S001].
Two 2D scans were performed for each crystal using a 137Cs source with the 1 mm diameter pinhole collimator with a $2\times 2$ mm2 pitch. One with the crystal placed horizontally and one vertically.
@@ 2.1 PSCS
The PSCS algorithm is an established technique for generating experimental bases from scanning table data. The PSCS algorithm relies on comparing signals from two intersecting scans (horizontal and vertical). The underlying principle is that signals originating from the same physical interaction point should have similar pulse shapes, regardless of the scan direction. A $\chi^{2}$ minimisation is employed to quantify the similarity between signals from the two scans using the following formula:
@@
where $V_{ch}(i)$ and $H_{ch}(i)$ represent the signal amplitudes of channel $ch$ at time sample $i$. The signals are digitised at a sampling frequency of 100 MHz, meaning each time sample corresponds to a 10 ns interval. Here, $\sigma$ represents the noise amplitude, and $N$ is the total number of samples compared (37 channels $\times$ 120 samples). Signal pairs with a $\chi^{2}$ value below a predefined threshold are considered to originate from the intersection point of the two scans. This process is iterative, refining the selection of signals to create a basis representative of the interaction position. However, the optimal $\chi^{2}$ threshold varies depending on the location within the crystal, requiring careful adjustment and complicating the automation of the process. This, combined with the inherently large number of comparisons required for each position, leads to significant computational cost and complexity in the implementation of the PSCS algorithm.
PSCS-derived bases for the S001 crystal [decan2020] were provided by collaborators at IPHC Strasbourg for comparison with the neural network-generated bases.
@@ 2.2 Data Preprocessing
Data acquisition was performed using TNT2 [TNT_digi] digital pulse processors. The resulting raw data underwent several pre-processing steps:
@@
    Conversion and Compression: Raw data was converted to the AGATA Data Format (ADF) [agapro] using the Scanning Table Data Processing (STDPro) software [scanneddatareader]. This step also compressed the data significantly ($\sim$ factor of 5).
@@
    Coincidence Filtering: To filter out random coincidences, events occurring in segments with low statistics were removed. A baseline was established by calculating the average event count from all segments not illuminated by the source. Any event whose fired segment had a total count less than three times this baseline was subsequently removed from the dataset.
@@
    Timing Alignment: Core and segment traces were time-aligned using standard AGATA workflow procedures [stez].
@@
    Trace Sampling: Super-traces were constructed and subsequently reduced in size. From the original 120 samples per trace, only the 60 samples corresponding to the rising edge and plateau were retained, as this window encapsulates the position-sensitive information, and it matches the number of samples used in the standard AGATA workflow. This yields a final super-trace dimension of $(36\text{ segments}+1\text{ core})\times 60\text{ samples}=37\times 60=2220$ features per event.
@@
    Format Conversion: Data was converted from ROOT format to NumPy arrays using the ‘uproot’ library [uproot] for compatibility with machine learning frameworks.
@@
    Dataset Preparation: The datasets from the vertical and horizontal scans were combined, and the event sequence was randomly permuted to ensure a homogeneous mixture of data points. The combined dataset was saved in HDF5 format which allows for loading specific events from the files without the need to load the entire file into memory, a feature that becomes extremely important when building the experimental basis.
@@
    Train/Validation Split: The full dataset for each crystal (around $10^{8}$ events per crystal) was split into a training set (90%) and a validation set (10%) to monitor model performance and prevent overfitting.
The training dataset for the NN model consisted of events within the energy ranges of 300-480 keV (upper Compton continuum) and 650-670 keV (photopeak region). This selection was based on a preliminary analysis indicating that these energy windows correspond to the lowest prediction errors achievable by the model. By concentrating the training on these high-fidelity regions, the aim was to optimise the model’s ability to learn the position-dependent features from the most informative signals, while reducing the impact of potentially less reliable data from lower energies or complex scattering regions known to exhibit higher errors and potentially more background contamination.
@@ 2.3.2 Network Architecture
The NN architecture employed is based on Long Short-Term Memory (LSTM) layers [LSTM]. As the detector signals (super-traces) represent time-sequential data where the temporal evolution encodes position information, LSTMs are particularly suitable. They are a type of Recurrent Neural Network (RNN) specifically designed to effectively learn and remember long-term dependencies within sequences. This capability is crucial for capturing the subtle features across the entire pulse shape relevant for accurate position determination. Moreover, LSTMs has been shown to be robust to time misalignment [FABIAN], which is crucial to successfully process AGATA signals. These LSTM layers are followed by dense layers for the final position regression. The input to the model is the super-trace, formatted as a single continuous vector of 2220 data points, the architecture includes an initial reshaping step to recover the temporal structure required by the LSTM. The detailed architecture, illustrated schematically in Fig. 2, consists of the following layers:
@@
Figure 2: Schematic diagram illustrating the NN architecture.
A systematic study of the network architecture indicated that increasing the number or size of dense layers did not improve accuracy, while the number of LSTM units affected convergence speed but not final accuracy if training iterations were adjusted proportionally.
@@ 2.3.3 Training Paradigm
The models were trained using the Adam optimiser [kingma2014adam] with a learning rate of $10^{-4}$. A batch size of 500 super-traces per iteration was used, with a total of $3\times 10^{4}$ iterations per model. The super-traces were normalised by scaling each one so that its peak amplitude equaled 1000 arbitrary units. This normalisation procedure relies on the standard assumption that the fundamental shape of a signal is independent of the total energy deposited. This approach is consistent with the established AGATA PSA methodology. The masked Euclidean distance (Eq. 2) served as the primary loss function. Throughout this work, the “error” is defined as the loss value, unless otherwise specified.
@@
  • Training on a uniformly sampled subset (equal number of super-traces per scanned point): This drastically reduced the dataset size (to  2% of the original) and resulted in poorer overall performance [jrjc].
@@
Figure 3 displays the prediction error calculated on the two known axes plotted against the segment index (0–35), where segments 0-5 correspond to Sector A (Layers 1-6), segments 6-11 correspond to Sector B (Layers 1-6), and so on, up to segments 30-35 for Sector F. Two distinct patterns are observable in the error distribution. First, the error is significantly higher for the rear segments (Layers 5 and 6) across all sectors. This is attributed to the larger physical volume of these segments, as well as reduced statistics caused by beam attenuation during the vertical scan (front entry). Second, a general increase in error is observed in Sectors E and F; this arises from attenuation during the horizontal scan. In this configuration, the detector was oriented with Sectors B and C facing the source. Consequently, the beam was attenuated by the crystal bulk before reaching the opposing sectors (E and F), resulting in lower statistics. Despite these physical variations, the per-segment training approach yielded the lowest average error and mitigated the distribution bias observed with a single global model. All subsequent results are based on this per-segment training strategy.
Figure 3: The error ($m$) on the validation dataset at photopeak energy versus segment number for different training methods: using a uniformly sampled subset (blue), using the full dataset with a single model trained with balanced batches per segment (orange), and training one model per segment (green). Error is calculated on the two known axes. The segments index is explained in the text.
@@ 2.3.4 Overfitting Mitigation
As the training of the model progressed, it started to exhibit signs of overfitting. Figure 4 tracks the evolution of the loss function throughout the training process, plotted against the number of iterations. The blue curve, representing the training loss, shows a continuous decline, indicating that the model is successfully minimising error on the data it is actively learning from. In contrast, the validation loss (orange curve), calculated on data the model has never seen, initially decreases but eventually reaches a minimum before beginning to rise. This divergence indicates that the model has started to memorise the specific noise of the training set rather than learning generalisable features, leading to degraded performance on new data.
To combat this overfitting and encourage generalisation, a Gaussian noise layer was added after the input layer in the network architecture. This layer adds random noise, drawn from a Gaussian distribution, to each sample of the input super-trace during training. Various noise levels (standard deviations) were experimentally tested during model development. While higher noise levels could further reduce the discrepancy between training and validation loss, they also tended to degrade the overall prediction accuracy (increase the minimum validation loss). Conversely, lower noise levels were less effective at controlling the overfitting. A standard deviation of 5 keV was ultimately selected as it provided the best observed balance between these competing factors. As shown in Fig. 4, the addition of this 5 keV noise significantly reduced the gap between training and validation loss (green and red curves, respectively) compared to training without noise, and more importantly, led to a lower converged validation loss, indicating improved model generalisation.